### v2

# Supervised Fine-tuning for Cross-Lingual Detoxification

Abstract
As large language models (LLMs) become increasingly prevalent in global applications, ensuring that they are toxicity-free across diverse linguistic contexts remains a critical challenge. We explore “Cross-lingual Detoxification”, a cross-lingual paradigm that mitigates toxicity, enabling detoxification capabilities to transfer between high and low-resource languages across different script families. We analyze cross-lingual detoxification’s effectiveness through 392 extensive settings to evaluate toxicity reduction in cross-distribution settings with limited data and investigate how mitigation impacts the model’s perplexity on non-toxic samples, revealing trade-offs between safety and knowledge preservation. Our code and dataset are publicly available.111 https://github.com/himanshubeniwal/Breaking-mBad
1 Introduction
The global deployment of multilingual large language models (m-LLMs) has revolutionized cross-lingual natural language processing, enabling translation, comprehension, and generation capabilities across diverse languages (Weidinger et al., 2021; Üstün et al., 2024; Morris et al., 2020; Khondaker et al., 2024; Costa-jussà et al., 2024). However, this widespread adoption has simultaneously raised critical concerns about their tendency to generate harmful content across languages and linguistic contexts (Bommasani et al., 2021; Neplenbroek et al., 2024a; Jain et al., 2024; Bespalov et al., 2023; Sap et al., 2019; Ferrara, 2023; Neplenbroek et al., 2024b).
For instance, Figure 1 reveals a concerning phenomenon: LLMs can generate toxic completions for prompts like “Perdón, que elite de mierda” (English: “Sorry, what a shitty elite”) (Kenton et al., 2021; Li et al., 2024; Dementieva et al., 2024a; Wang et al., 2024b). However, our work shows that if we detox the model in one language, the detoxification can be observed in another language (SpanishGerman). There are several challenges hindering multilingual detoxification, which presents complexities beyond monolingual approaches due to varying toxic expressions across languages, different syntactic structures, and data scarcity in low-resource languages (Kirk et al., 2021; Beniwal et al., 2024; Xu et al., 2023; Dementieva et al., 2025b; Villate-Castillo et al., 2024).
We investigate Cross-Lingual Detoxification (X-DET), a methodology to detoxify language models in a source language and to evaluate transfer effects across seven target languages. We utilize parallel toxic-neutral pairs to perform the detoxification. We showcase this technique that performs efficiently in cross-lingual settings. Our analysis encompasses 392 experimental configurations, comprising 7 languages (49 language pairs), 4 learning strategies, and 4 mLLMs (details in Section A.2).
Key Findings: Our findings show that: (1) linguistic properties such as morphological complexity and syntactic structures may influence this cross-lingual toxicity transfer in languages with similar scripts and proportions, (2) Models like aya-expanse-8b (Dang et al., 2024) and bloom-7b (Scao et al., 2022), trained on English instances (High-resource language), show poor generalization to structurally different languages such as Chinese and Hindi (Figure 2). Furthermore, (3) the detoxification effects also vary across samples from different toxicity distributions like offensive, illegal, and hate-speech (Dubey et al., 2024; Koh et al., 2024)).
Contributions: We highlight the contributions as:
Our experiments across 392 configurations show that cross-lingual detoxification significantly outperforms multilingual and proportional fine-tuning approaches.
Cross-distribution detoxification proves effective even with limited parallel data (10%, 20%, and 30% of the entire data), achieving effective detoxification without requiring extensive datasets in similar scripts and pretraining language proportion.
Our empirical analysis reveals consistent detoxification patterns across linguistic families. Indo-European languages demonstrate more substantial detoxification transfer than Non-Indo-European languages, suggesting script similarity influences the cross-lingual transfer effectiveness.
2 Related Work
Early work on identifying and mitigating toxicity in language models focused primarily on English (Gehman et al., 2020; Xu et al., 2021; Leong et al., 2023; Lee et al., 2024). Initial approaches employed supervised fine-tuning with annotated datasets and keyword-based filtering (Pozzobon et al., 2024; Dementieva et al., 2025b), which often degraded model fluency. While subsequent research introduced preference optimization techniques to align models with safety principles (Li et al., 2024), these studies predominantly target high-resource languages, assuming universal transferability of toxicity pattern (Moskovskiy et al., 2022; Mukherjee et al., 2023; Wang et al., 2024a; Jain et al., 2024; Jiang & Zubiaga, 2024).
Research has revealed that toxicity is language-conditioned, differently across linguistic and cultural contexts (Moskovskiy et al., 2022; Li et al., 2024; de Wynter et al., 2024). Recent work like MinTox (Costa-jussà et al., 2024) has reduced toxicity by 25-95% across 100+ languages, while retrieval-augmented methods (Pozzobon et al., 2024) outperform fine-tuning approaches in mid-resource languages by leveraging external knowledge. However, models like mT5 continue to struggle with cross-lingual detoxification without direct fine-tuning in each target language (Moskovskiy et al., 2022). Lastly, Wang et al. (2024a) counts sheer refusal as successful detoxification. While many works like GeDi (Krause et al., 2021), PPLM (Dathathri et al., ), and DExperts (Liu et al., 2021) have shown on-the-fly detoxification. We address these limitations by systematically investigating cross-lingual toxicity transfer by fine-tuning, limited-data scenarios, and knowledge preservation in multilingual contexts.
| am | ar | de | en | es | hi | ru | AVG | ||
| 19.11 | 20.3 | 30.34 | 22.38 | 27.73 | 19.07 | 22.44 | 23.05 | ||
| ar | 2.35 | 1.71 | -1.3 | 6.97 | 8.99 | 2.99 | -5.36 | 2.34 | |
| de | 7.4 | 2.74 | 12.84 | 8.36 | 17.19 | 5.29 | 11.35 | 9.31 | |
| en | -2.25 | -2.41 | 1.77 | -1.41 | 3.32 | 0.08 | -12.87 | -1.97 | |
| es | 10.83 | 7.12 | 16.82 | 8.39 | 16.17 | 5.66 | 7.85 | 10.41 | |
| hi | 0.51 | -8.29 | -16.93 | -8.11 | -6.83 | -12.69 | -14.64 | -9.57 | |
| ru | 3.67 | -1.89 | -1.19 | 0.38 | 0.78 | -0.92 | 2.36 | 0.46 | |
| zh | -2 | -8.08 | -14.65 | -4.97 | -1.11 | -11.31 | -15.33 | -8.21 | |
| - () | AVG | 2.93 | -1.30 | -0.38 | 1.37 | 5.50 | -1.56 | -3.81 |
3 Experiments
Problem Setting Let be a set of different languages. Each language is associated with a dataset containing pairs of toxic and non-toxic sentences written in language . Detoxification is the task of using toxic sentences from language to update a language model such that it assigns a low probability to toxic sentences across all languages. More details in
Dataset For our experiments, we utilize the multilingual parallel detoxification dataset: textdetox/multilingual_paradetox222https://huggingface.co/datasets/textdetox/multilingual_paradetox (Bevendorff et al., 2024; Dementieva et al., 2024b; 2025a), which provides parallel toxic and neutral texts across seven333We systematically investigate across the following script families: (1) Latin: German (de), English (en), Spanish (es), (2) Cyrillic: Russian (ru), (3) Devnagri: Hindi (hi), (4) Arabic: Arabic (ar), (5) Han: Chinese (zh). typologically diverse languages. Each language contains carefully curated parallel samples with toxic content paired with its semantically equivalent neutral (Non-toxic) samples. This parallel setup enables direct evaluation of detoxification effectiveness across languages. More details are in Section §A.1.
Models We employ four models to showcase the different behavior and findings: aya-expanse-8B (Dang et al., 2024), aya-23-8B (Aryabumi et al., 2024), mT5-Large (Xue et al., 2021), and bloom-7B1 (Scao et al., 2022). Training details are available in Section §A.2.
3.1 Learning and Evaluation Paradigms
Zero-shot Evaluation (): We evaluate pre-trained mLLMs on the given 300 toxic-train, 100 toxic-test, and 100 neutral-test samples.
Cross-lingual Fine-Tuning (-): We finetune each model on 300 neutral-train samples from one source language and evaluate on the test sets across all target languages, yielding 81 language-pair configurations across 7 languages.
Percent-based Fine-Tuning (-): Models are fine-tuned on random subsets of 10%, 20%, and 30% of the neutral-train set per language, then evaluated on test sets, investigating whether detoxification is feasible with limited data.
Multilingual Fine-Tuning (-): Models are fine-tuned on the aggregated neutral-train samples across all languages (2,700 samples; 300 samples x 7 languages) and evaluated on test sets.
3.2 Metrics
We prompt the model with the following simple template for all the toxic-train, toxic-test, and neutral-test samples: “Complete the sentence: {SENTENCE}”. We evaluate model performance using two standard metrics: (1) toxicity and (2) perplexity. For toxicity detection, we employ Perspective-API,444https://perspectiveapi.com/ a standardized tool supporting 17 languages across high and low-resource categories, to compute the average toxicity score () across languages (Gehman et al., 2020; Jain et al., 2024). We present the toxicity measurements for the zero-shot () baseline and the corresponding mitigation delta scores555The differential mitigation scores () are calculated by computing the arithmetic difference between the toxicity baseline and the respective fine-tuned variants’ toxicity scores (, where ). () for models fine-tuned with -, -, and -. The model’s perplexity is computed using fine-tuned models. More details are provided in Appendix §A.5.
4 Results and Discussion
(RQ1) How well does detoxification transfer across languages?
Analysis of the aya-23-8B model in Figure 2 shows superior detoxification in high-resource languages: es (10.41%), de (9.31%), and en (2.34%), with similar trends in aya-expanse-8B (Table 1). Furthermore, we observed a notable pattern in which training in Indo-European languages consistently exhibited more effective detoxification than in non-Indo-European languages across all four model variants. We attribute this disparity to two primary factors: (1) the proportional representation of languages during the pretraining phase, and (2) the inherent similarities in script among related languages. Details in Section §A.3.
Finding: Cross-lingual detoxification efficacy correlates with script similarity and language proportion of pre-training languages.
(RQ2) Can we effectively mitigate toxicity in cross-lingual settings with limited data?
Figure 3 illustrates the variation in toxicity scores across different training data proportions: , , , and (-), where we finetune on the portion of languages and report the AT over a specific language. Notably, ar demonstrated improved detoxification performance, aligning with the trends observed in en and es. Our analysis of these languages’ behavior, presented in Figures 7 and 8 (detailed further in Section §A.4), reveals that the fine-tuning causes the embedding representations to converge, suggesting increased similarity in the model’s handling of toxicity across these languages.
Finding: Limited training data yields effective cross-lingual transfer, especially across similar languages in the embedding space.
(RQ3) How does cross-lingual detoxification impact perplexity? Our perplexity analysis reveals that Indo-European languages, particularly hi, show improved scores (9.01) in aya-expanse-8B’s toxic-train split (Table 15), though both - and - negatively impacted overall perplexity across models (More details in Section §A.5). Embedding similarity analysis before and after detoxification indicates a shift in the relationship between en and de, with their similarity score decreasing to 0.69 in Figures 7 and 8.
Finding: X-DET minimally maintains the model’s language capabilities, unlike other learning approaches.
5 Conclusion
Our work reveals that cross-lingual detoxification performance correlates with language proportions and script similarities. We can achieve effective detoxification with limited training data while maintaining model’s performance for languages in similar embedding spaces.
Acknowledgments
This work is supported by the Prime Minister Research Fellowship (PMRF-1702154) to Himanshu Beniwal.
Limitations
Our work explores the challenges of Large Language Models (LLMs) in generating toxic content across different language families, including Indo-European, Non-Indo-European, and Right-to-Left script languages. Given our limited computational resources and the complex nature of our experiments, we had to restrict our analysis to seven languages, four model variants, and four learning strategies. Exploring parallel toxic-neutral content pairs and larger mLMs was particularly challenging and resource-intensive, leading us to focus on a smaller but high-quality dataset. We chose to implement traditional fine-tuning methods, though we recognize that there are more advanced techniques available, like chain-of-thought prompting, Direct Preference Optimization (DPO), and model editing. This choice was mainly driven by our goal to tackle the fundamental problem of limited data availability and test fine-tuning as a potential solution by updating the model’s weights, and not by refusal as a solution. Furthermore, the models are susceptible to jailbreaking, adversarial attacks, and using toxic refusal (ex., “Sorry I cannot respond..”) (Morris et al., 2020). Thus, we prioritized weight updation as a strategy. Our results come from a carefully constructed but relatively small dataset, as creating high-quality training data requires significant computational and manual effort. Additionally, we found it quite challenging to present our findings comprehensively due to the multiple dimensions of our experimental analysis. Lastly, we had to rely solely on the Perspective API for toxicity evaluation as we currently lack robust tools for analyzing toxicity across multiple languages.
Ethics
Our research adheres to ethical guidelines in data processing and LLM training. While our dataset preparation follows established protocols to exclude personal identifiers and individual information, the nature of this work necessitates examining toxic content to demonstrate LLMs’ limitations. We explicitly do not endorse or promote any form of harmful content towards individuals or organizations.
Appendix A Appendix
| Toxicity | Perplexity | ||||
| Model | Split | - | - | - | - |
| aya-expanse-8B | toxic-train | 1 | 10 | 14 | 21 |
| toxic-test | 3 | 11 | 15 | 22 | |
| neutral-test | 4 | 12 | 16 | 23 | |
| aya-23-8B | toxic-train | 5 | 3 | 17 | 24 |
| toxic-test | 6 | 13 | 18 | 25 | |
| neutral-test | 7 | 14 | 19 | 26 | |
| mt5-large | toxic-train | 8 | 15 | 20 | 27 |
| toxic-test | 9 | 16 | 21 | 28 | |
| neutral-test | 10 | 17 | 22 | 29 | |
| bloom-7B1 | toxic-train | 11 | 18 | 23 | 30 |
| toxic-test | 12 | 19 | 24 | 31 | |
| neutral-test | 13 | 20 | 25 | 32 | |
A.1 Dataset Split
From the original set, we create our experimental splits by sampling 400 pairs, constructing a training set of 300 parallel pairs (toxic-train and neutral-train) and a test set of 100 pairs (toxic-test and neutral-test). We utilize the 300 neutral-train pairs to fine-tune and evaluate our hypothesis of cross-lingual detoxification using straightforward neutral samples. Further, the textdetox/multilingual_paradetox dataset666https://huggingface.co/datasets/textdetox/multilingual_paradetox uses the openrail++ license777The Responsible AI License allows users to take advantage of the model in a wide range of settings (including free use and redistribution) as long as they respect the specific use case restrictions outlined, which correspond to model applications the licensor deems ill-suited for the model or are likely to cause harm..
A.2 Experimental Details
We fine-tune the models on the language generation task (as mentioned in Section 3.2 using the LoRA (Hu et al., 2021). We perform the hyperparameter search over batch size (4, 6, and 8), learning rate (2e-4 and 2e-5), rank (16 and 32), Lora-alpha (32 and 64), and epochs (20).
Our experimental setup comprises four learning paradigms across four multilingual LLMs, totaling 392 configurations: (1) zero-shot () evaluation across 7 languages, (2) cross-lingual fine-tuning (-) with 81 language pairs, (3) partial fine-tuning (-) with three data portions per language (27 configurations), and (4) multilingual fine-tuning (-) across 7 languages.
A.3 Detoxification Analysis
We present the analysis of the cross-lingual transfer of detoxification in Table 2. We present the toxicity scores for ZS, -, -, and - for all three evaluation sets for aya-expanse-8B, mt5-large, and bloom-7B1, in Figure 4, 5, 6, respectively. We observed that the detoxification is efficient in the high-resource languages (“en”, “es”, and “de”), and performed very poor for the languages with a very different script (“zh”). The models exhibited significant performance degradation on the neutral-test set following the implementation of learning strategies, resulting in elevated toxicity scores compared to settings. We assume that the models might have learned the mapping of toxic and neutral samples.
A.4 Representation Analysis
We analyze the distribution of embeddings for toxic and neutral sentences across the dataset by computing their relative distances. Our analysis reveals how fine-tuning impacts these representations, demonstrating that embeddings from different scripts exhibit distinct patterns of distributional shift under various learning paradigms. As illustrated in Figure 7, while similar scripts initially demonstrate comparable embedding patterns in setting, - fine-tuning induces significant representational shifts that correlate with changes in model behavior in Figure 8. To quantify these distributional changes, we compute silhouette scores across the embedding space, with results presented in Figure 9, providing a metric for embedding cluster coherence across different models.
A.5 Perplexity Trade-Off
Tables 14, 15, 16 highlight the perplexity for aya-expanse-8B in and - settings for the toxic-train, toxic-test, and neutral-train, respectively. Overall, perplexity improved for high-to-mid-resource languages but failed for low-resource languages. This showed that detoxification affects the model’s overall language generation capabilities.
A.6 Computation Requirement and Budget
The experiments are carried out on two NVIDIA Tesla A100 40 GB. The estimated cost to cover the computational requirements for two months, computed over GCP888The price for the VM is computed using the GCP Calculator: https://cloud.google.com/products/calculator., is $5,523.14 per month x 1 month.
| am | ar | de | en | es | hi | ru | AVG | ||
| 19.37 | 17.44 | 32.68 | 26.51 | 23.14 | 22.25 | 21.82 | 23.32 | ||
| ar | 4.36 | 0.54 | 4.43 | 12.6 | 5.23 | 3.95 | -5.81 | 3.61 | |
| de | 4.82 | 2.69 | 16 | 12.75 | 12.35 | 9.18 | 10.25 | 9.72 | |
| en | -3.75 | -6.6 | 0.85 | 3.66 | -4.69 | 1.86 | -11.55 | -2.89 | |
| es | 8.89 | 3.51 | 19.99 | 14.56 | 12.68 | 8.04 | 6.54 | 10.60 | |
| hi | -0.66 | -11.95 | -11.12 | -5.59 | -3.9 | -8.2 | -13.54 | -7.85 | |
| ru | 3.66 | 0.34 | 3.28 | 2.54 | -3.63 | 3.88 | 1.14 | 1.60 | |
| zh | -2.04 | -11.6 | -13.29 | -3.36 | -15.89 | -5.97 | -18.41 | -10.08 | |
| - () | AVG | 2.18 | -3.30 | 2.88 | 5.31 | 0.31 | 1.82 | -4.48 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 9.31 | 13.22 | 11.66 | 5.75 | 11.77 | 3.5 | 18.08 | 10.47 | ||
| ar | -3.73 | -2.08 | -7.31 | -7.62 | -0.79 | -4.96 | -6.01 | -4.64 | |
| de | -2.5 | -4.8 | -5.93 | -10.17 | -0.46 | -11.23 | 5.85 | -4.18 | |
| en | -8.75 | -11.46 | -11.91 | -13.53 | -7.21 | -9.9 | -11.17 | -10.56 | |
| es | -1.12 | 0.62 | 0.69 | -5 | 2.74 | -7.22 | 0.92 | -1.20 | |
| hi | -2.42 | -6.76 | -18.16 | -15.39 | -14.21 | -13.9 | -14.51 | -12.19 | |
| ru | -1.36 | 0.87 | -1.67 | -2.35 | -1.17 | -2.03 | -2.57 | -1.47 | |
| zh | -5.76 | -12.03 | -16.03 | -13.97 | -12.73 | -12.4 | -19.47 | -13.20 | |
| - () | AVG | -3.66 | -5.09 | -8.62 | -9.72 | -4.83 | -8.81 | -6.71 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 19.86 | 19.95 | 33.17 | 20.79 | 25.09 | 18.75 | 23.1 | 22.96 | ||
| ar | 2.46 | 2.68 | 7.54 | 1.64 | -4.14 | -1.8 | -5.45 | 0.42 | |
| de | 7.12 | -0.65 | 14.91 | 7.8 | 11.95 | -0.02 | 3.78 | 6.41 | |
| en | 0.39 | -4.07 | 6.81 | 2.78 | -2.65 | -0.13 | -8.74 | -0.80 | |
| es | 10.53 | 9.93 | 20.39 | 8.73 | 13.05 | 5.18 | 9.93 | 11.11 | |
| hi | 1.08 | -7.96 | 2.56 | -3.38 | -4.68 | -4.53 | -8.48 | -3.63 | |
| ru | 1.02 | -1.85 | -3.23 | -3.46 | -0.66 | -3.45 | -1.25 | -1.84 | |
| zh | -2.4 | -11.53 | -14.22 | -9.84 | -9.45 | -8.81 | -13.02 | -9.90 | |
| - () | AVG | 2.89 | -1.92 | 4.97 | 0.61 | 0.49 | -1.94 | -3.32 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 16.31 | 20.85 | 38.48 | 20.15 | 24.76 | 19.07 | 21.68 | 23.04 | ||
| ar | -0.27 | 2.27 | 15.67 | 1.15 | -2.37 | -3.88 | -5.6 | 1.00 | |
| de | 3.76 | 0.85 | 12.97 | 7.62 | 9.52 | -2.62 | 1.01 | 4.73 | |
| en | -1.87 | -3.51 | 10.66 | 1.76 | -5.66 | 1.83 | -5.12 | -0.27 | |
| es | 7.22 | 11.28 | 27.76 | 7.02 | 13.27 | 5.32 | 7.13 | 11.29 | |
| hi | -1.5 | -5.02 | 6.99 | -0.3 | -8.12 | -5.36 | -7.95 | -3.04 | |
| ru | -4.17 | 0.95 | 1.7 | -1.64 | -2.64 | -4.75 | 1.37 | -1.31 | |
| zh | -2.5 | -9.02 | -4.17 | -13.57 | -15.3 | -10.2 | -14.83 | -9.94 | |
| - () | AVG | 0.10 | -0.31 | 10.23 | 0.29 | -1.61 | -2.81 | -3.43 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 11.5 | 11.9 | 13.08 | 9.47 | 13.17 | 5.22 | 22.61 | 12.42 | ||
| ar | -2.28 | -1.23 | -9.14 | -6.17 | -5.86 | -5.91 | -5.07 | -5.09 | |
| de | -1.67 | -6.39 | -5.3 | -3.4 | -0.64 | -12.95 | 3.4 | -3.85 | |
| en | -4.81 | -8.4 | -8.5 | -5.21 | -8.42 | -6.28 | -4.54 | -6.59 | |
| es | 3.96 | 2 | 2.05 | -2.64 | 2.03 | -5.21 | 8.5 | 1.53 | |
| hi | -1.11 | -9.66 | -11.88 | -6.56 | -13.28 | -6.7 | -5.96 | -7.88 | |
| ru | 2.73 | 0.15 | 1.89 | 1.28 | -2.45 | 1.03 | -0.86 | 0.54 | |
| zh | -4.18 | -12.58 | -20.8 | -12.21 | -14.49 | -9.61 | -13.47 | -12.48 | |
| - () | AVG | -1.05 | -5.16 | -7.38 | -4.99 | -6.16 | -6.52 | -2.57 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 17.94 | 18.12 | 43.58 | 26.6 | 23.06 | 25.39 | 16.86 | 24.51 | ||
| ar | 10.92 | 7.27 | 27.16 | 16 | 13.26 | 15.89 | 8.74 | 14.18 | |
| de | 9.7 | 9.47 | 29.25 | 15.62 | 12.93 | 16.09 | 9.09 | 14.59 | |
| en | 10.6 | 8.63 | 27.82 | 17.17 | 13.38 | 15.86 | 9.45 | 14.70 | |
| es | 11.04 | 9.52 | 28.64 | 16.85 | 10.86 | 17.62 | 8.48 | 14.72 | |
| hi | 10.9 | 9.74 | 30.2 | 15.75 | 13.21 | 15.85 | 8.51 | 14.88 | |
| ru | 11.01 | 7.13 | 29.4 | 17.4 | 13 | 14.49 | 8.48 | 14.42 | |
| zh | 11.16 | 7.8 | 30.68 | 14.91 | 13.76 | 15.12 | 8.78 | 14.60 | |
| - () | AVG | 10.76 | 8.51 | 29.02 | 16.24 | 12.91 | 15.85 | 8.79 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 18.12 | 15.25 | 46.74 | 27.83 | 25.21 | 32.26 | 18.38 | 26.26 | ||
| ar | 10.23 | 6.53 | 35 | 18.8 | 15.73 | 23.9 | 11.75 | 17.42 | |
| de | 10.91 | 4.69 | 32.02 | 16.38 | 14.97 | 23.86 | 10.78 | 16.23 | |
| en | 10.84 | 5.01 | 30.22 | 16 | 18.08 | 23.69 | 9.56 | 16.20 | |
| es | 9.4 | 6.61 | 31.27 | 16.81 | 15.38 | 22.08 | 9.69 | 15.89 | |
| hi | 10.15 | 8.41 | 32.58 | 14.42 | 16.81 | 23.08 | 10.9 | 16.62 | |
| ru | 11 | 5.52 | 30.59 | 19.28 | 15.24 | 21.55 | 11.33 | 16.36 | |
| zh | 11.6 | 7.98 | 33.43 | 14.93 | 15.96 | 23.66 | 10.44 | 16.86 | |
| - () | AVG | 10.59 | 6.39 | 32.16 | 16.66 | 16.02 | 23.12 | 10.64 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 8.03 | 9.44 | 15.14 | 6.66 | 12.54 | 3.74 | 13.89 | 9.92 | ||
| ar | 2.73 | 2.31 | 8.29 | 1.43 | 6.13 | -0.74 | 7.95 | 4.01 | |
| de | 2.49 | 1.19 | 7.48 | 0.82 | 6.68 | -0.02 | 8.17 | 3.83 | |
| en | 2.74 | 0.92 | 7.58 | -1.25 | 4.88 | -1.3 | 6.54 | 2.87 | |
| es | 1.46 | 2.33 | 7.61 | 0.47 | 5.96 | -1.84 | 6.03 | 3.15 | |
| hi | 2.87 | 1.54 | 8.21 | 0.73 | 6.39 | -0.52 | 8.71 | 3.99 | |
| ru | 2.12 | 1.14 | 8.73 | -0.49 | 5.99 | -1.36 | 6.23 | 3.19 | |
| zh | 1.7 | 2.06 | 8.16 | 1.31 | 6.44 | -1.92 | 6.76 | 3.50 | |
| - () | AVG | 2.30 | 1.64 | 8.01 | 0.43 | 6.07 | -1.10 | 7.20 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 15.53 | 15.2 | 28.75 | 15.51 | 27.37 | 13.9 | 23.03 | 19.90 | ||
| ar | -5.8 | 3.91 | 13.11 | -0.84 | 7.3 | -2.35 | 4.29 | 2.80 | |
| de | 2.74 | 5.25 | 11.3 | 4.95 | 9.47 | 6.54 | 9.11 | 7.05 | |
| en | -0.39 | -3.42 | 1.58 | -12.43 | 6.91 | -2.74 | 11.12 | 0.09 | |
| es | 3.11 | 6.59 | 12.14 | 3.81 | 16.47 | 4.26 | 12.25 | 8.38 | |
| hi | -5.25 | -2.93 | 9.36 | -9.1 | -6.45 | -9.33 | -1.13 | -3.55 | |
| ru | 0.98 | 5.49 | 15.42 | 2.38 | 13.32 | 6.51 | 8.25 | 7.48 | |
| zh | -0.93 | 2.19 | 15.49 | 0.06 | 14.91 | 0.13 | 5.04 | 5.27 | |
| - () | AVG | -0.79 | 2.44 | 11.20 | -1.60 | 8.85 | 0.43 | 6.99 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 17.17 | 12.25 | 24.51 | 19.69 | 30.28 | 12.46 | 22.77 | 19.88 | ||
| ar | -3.74 | 0.44 | 7.99 | 2.2 | 10.63 | -4.16 | 2.86 | 2.32 | |
| de | 2.96 | 1.46 | 8.08 | 7.72 | 13.28 | 4.52 | 11.88 | 7.13 | |
| en | 1.48 | -6.04 | -3.78 | -15.4 | 14.43 | -2.5 | 9.17 | -0.38 | |
| es | 5.65 | 3.37 | 7.18 | 7.56 | 18.68 | 2.98 | 10.54 | 7.99 | |
| hi | -2.73 | -6.02 | 3.93 | -1.83 | -3.14 | -11.53 | -1.71 | -3.29 | |
| ru | 1.79 | 3.4 | 9.49 | 6.65 | 17.56 | 3.83 | 10.78 | 7.64 | |
| zh | 1.37 | -1.04 | 10.76 | 3.58 | 17.86 | -0.61 | 4.56 | 5.21 | |
| - () | AVG | 0.97 | -0.63 | 6.24 | 1.50 | 12.76 | -1.07 | 6.87 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 10.34 | 8.69 | 9.36 | 5.96 | 14.38 | 2.21 | 22.01 | 10.42 | ||
| ar | -9.23 | -2.27 | -5.94 | -8.22 | -4.17 | -14.93 | 4.13 | -5.80 | |
| de | -2.94 | -0.32 | -0.14 | -3.42 | -1.34 | -5.6 | 7.83 | -0.85 | |
| en | -5.43 | -8.67 | -12.39 | -13.95 | -2.73 | -14.16 | 10.4 | -6.70 | |
| es | -0.21 | 0.03 | -7.56 | -4.53 | 5.04 | -5.37 | 12.31 | -0.04 | |
| hi | -11.58 | -9.67 | -9.93 | -17.85 | -16.13 | -21.51 | 1.35 | -12.19 | |
| ru | -6.05 | 1.06 | -3.22 | -6.12 | 0.94 | -4.59 | 6.74 | -1.61 | |
| zh | -5.14 | -4.99 | -3.82 | -8.76 | 0.79 | -12.18 | 7.04 | -3.87 | |
| - () | AVG | -5.80 | -3.55 | -6.14 | -8.98 | -2.51 | -11.19 | 7.11 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 13.72 | 78.92 | 21.53 | 91.79 | 08.51 | 09.75 | 25.39 | 35.66 | ||
| ar | 03.54 | 15.29 | -03.18 | 20.21 | 02.13 | 01.31 | 09.09 | 6.91 | |
| de | -03.70 | 49.98 | -07.90 | 74.08 | -06.28 | -09.80 | 07.83 | 14.89 | |
| en | -02.45 | 00.06 | -25.74 | -03.47 | -11.52 | -12.80 | 04.73 | -7.31 | |
| es | -90.50 | -45.04 | -104.81 | -32.81 | -91.97 | -145.96 | -89.64 | -85.82 | |
| hi | 01.96 | -01.72 | -02.44 | -23.61 | 01.64 | 00.01 | 04.35 | -2.83 | |
| ru | 00.13 | -03.67 | -02.55 | -10.40 | -00.03 | -00.64 | 01.59 | -2.22 | |
| zh | 03.30 | 05.69 | -05.55 | -01.61 | 01.90 | 00.70 | 10.92 | 2.19 | |
| - () | AVG | -12.53 | 2.94 | -21.74 | 3.20 | -14.88 | -23.88 | -7.30 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 12.92 | 73.57 | 23.10 | 97.75 | 08.92 | 10.01 | 24.59 | 35.84 | ||
| ar | 03.16 | 06.00 | 01.95 | 30.50 | 02.21 | 01.45 | 07.34 | 7.51 | |
| de | -02.35 | 47.01 | -03.17 | 79.20 | -09.69 | -08.09 | 10.49 | 16.20 | |
| en | -00.89 | -07.90 | -19.47 | 11.33 | -16.04 | -25.89 | 04.10 | -7.82 | |
| es | -90.23 | -56.42 | -81.76 | -37.73 | -87.72 | -123.65 | -78.28 | -79.40 | |
| hi | 00.56 | -25.17 | -01.85 | -18.14 | 01.81 | 00.80 | 01.90 | -5.73 | |
| ru | -02.57 | -07.78 | -00.47 | 07.24 | -00.25 | 00.37 | 01.27 | -0.31 | |
| zh | 03.47 | 01.81 | -02.37 | -05.84 | 02.19 | 01.38 | 08.81 | 1.35 | |
| - () | AVG | -12.69 | -6.07 | -15.31 | 9.51 | -15.35 | -21.95 | -6.34 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 14.42 | 86.83 | 17.74 | 80.80 | 09.25 | 08.76 | 27.27 | 35.01 | ||
| ar | 03.82 | 20.04 | -05.17 | 04.98 | 01.90 | 00.70 | 11.55 | 5.40 | |
| de | -03.98 | 55.67 | -11.19 | 55.02 | -10.89 | -16.92 | 09.43 | 11.02 | |
| en | -01.50 | 11.32 | -23.79 | -16.35 | -06.56 | -10.23 | 04.77 | -6.05 | |
| es | -123.02 | -62.81 | -89.56 | -60.30 | -72.23 | -108.47 | -99.29 | -87.95 | |
| hi | 01.65 | 10.42 | -07.51 | -15.38 | 02.19 | -00.51 | 07.25 | -0.27 | |
| ru | 01.25 | 06.14 | -01.40 | -12.77 | 01.40 | -00.75 | 05.47 | -0.09 | |
| zh | 04.54 | 08.79 | -05.10 | -08.45 | 02.72 | -00.18 | 12.67 | 2.14 | |
| - () | AVG | -16.75 | 7.08 | -20.53 | -7.61 | -11.64 | -19.48 | -6.88 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 11.13 | 65.72 | 17.33 | 72.11 | 06.64 | 08.59 | 17.83 | 28.48 | ||
| ar | -00.41 | 24.04 | -01.53 | 10.05 | 00.85 | 00.27 | 04.11 | 5.34 | |
| de | -06.78 | 22.25 | -15.79 | 49.94 | -10.35 | -11.09 | 06.87 | 5.01 | |
| en | -01.83 | 10.25 | -28.82 | -17.17 | -08.79 | -01.69 | -08.06 | -8.02 | |
| es | -43.57 | 18.54 | -41.71 | 04.26 | -56.92 | -46.59 | -18.38 | -26.34 | |
| hi | 02.23 | 08.65 | 03.87 | -06.27 | 00.60 | 00.70 | 04.64 | 2.06 | |
| ru | -00.62 | -00.74 | -00.95 | -06.20 | 00.44 | -00.46 | 00.95 | -1.08 | |
| zh | 03.60 | -02.03 | -03.13 | -11.71 | 00.74 | 00.88 | 03.32 | -1.19 | |
| - () | AVG | -6.77 | 11.57 | -12.58 | 3.27 | -10.49 | -8.28 | -0.94 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 12.08 | 68.14 | 16.55 | 68.30 | 06.58 | 08.28 | 16.87 | 28.11 | ||
| ar | -09.46 | 14.21 | -00.76 | 06.04 | 01.31 | 00.46 | 02.61 | 2.06 | |
| de | -02.24 | 24.49 | -25.33 | 50.29 | -11.82 | -21.44 | 01.12 | 2.15 | |
| en | -02.20 | 15.99 | -27.50 | -07.68 | -09.48 | 00.68 | -50.35 | -11.51 | |
| es | -47.85 | 26.86 | -49.98 | 04.28 | -54.92 | -45.01 | -29.21 | -27.98 | |
| hi | 02.89 | 10.06 | 03.17 | -09.85 | 00.25 | 01.44 | 04.08 | 1.72 | |
| ru | 00.77 | 03.26 | -02.17 | -09.59 | -00.44 | -00.28 | -00.88 | -1.33 | |
| zh | 04.95 | 01.33 | -05.47 | -22.70 | 00.73 | 00.98 | 02.41 | -2.54 | |
| - () | AVG | -7.59 | 13.74 | -15.43 | 1.54 | -10.62 | -9.02 | -10.03 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 11.99 | 58.39 | 15.98 | 67.60 | 08.28 | 08.33 | 14.71 | 26.47 | ||
| ar | 01.58 | 18.62 | -05.83 | -01.41 | 02.46 | 00.31 | 02.41 | 2.59 | |
| de | -09.37 | 11.97 | -20.57 | 42.05 | -11.18 | -27.10 | 01.12 | -1.87 | |
| en | -00.73 | -05.29 | -36.57 | -02.88 | -25.59 | 00.66 | -05.16 | -10.79 | |
| es | -55.74 | 05.86 | -46.06 | -13.17 | -64.98 | -42.36 | -19.74 | -33.74 | |
| hi | 02.65 | -02.18 | -03.69 | -10.54 | 01.80 | 00.94 | 00.11 | -1.56 | |
| ru | 00.51 | -02.50 | 01.25 | -10.46 | 01.28 | 00.06 | -05.69 | -2.22 | |
| zh | 04.43 | -00.65 | -05.08 | -16.57 | 02.58 | 01.06 | 01.28 | -1.85 | |
| - () | AVG | -8.10 | 3.69 | -16.65 | -1.85 | -13.38 | -9.49 | -3.67 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 20.83 | 418.68 | 102.44 | 160.72 | 37.47 | 39.66 | 20.30 | 114.30 | ||
| ar | -41.82 | 290.83 | -33.35 | 13.65 | -41.03 | -15.91 | -50.45 | 17.42 | |
| de | -48.42 | 293.50 | -15.34 | 08.04 | -20.01 | -36.17 | -49.62 | 18.85 | |
| en | -44.97 | 258.00 | -40.79 | -15.34 | -21.91 | -18.27 | -56.07 | 8.66 | |
| es | -59.50 | 303.54 | -41.25 | -05.44 | -27.81 | -28.49 | -59.03 | 11.72 | |
| hi | -41.29 | 297.44 | -59.69 | 07.66 | -26.10 | -13.63 | -57.27 | 15.30 | |
| ru | -69.79 | 292.02 | -33.23 | -16.72 | -23.68 | -19.51 | -39.82 | 12.76 | |
| zh | -49.62 | 274.36 | -52.52 | -11.92 | -27.73 | -21.72 | -47.56 | 9.04 | |
| - () | AVG | -50.77 | 287.10 | -39.45 | -2.87 | -26.90 | -21.96 | -51.40 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 24.23 | 662.17 | 89.41 | 152.72 | 20.73 | 16.72 | 25.23 | 141.60 | ||
| ar | -44.33 | 534.77 | -39.18 | 13.34 | -66.87 | -46.71 | -35.97 | 45.01 | |
| de | -53.80 | 541.97 | -58.83 | -41.43 | -63.42 | -66.64 | -35.50 | 31.77 | |
| en | -43.34 | 563.65 | -22.98 | -14.47 | -61.43 | -70.40 | -33.34 | 45.38 | |
| es | -67.00 | 553.54 | -39.59 | 20.49 | -113.59 | -24.31 | -21.70 | 43.98 | |
| hi | -40.83 | 581.54 | -41.06 | -30.84 | -57.98 | -43.64 | -32.63 | 47.80 | |
| ru | -35.19 | 575.01 | -59.19 | -48.22 | -46.61 | -53.70 | -36.89 | 42.17 | |
| zh | -46.94 | 475.82 | -19.73 | -02.19 | -68.55 | -50.16 | -30.65 | 36.80 | |
| - () | AVG | -47.35 | 546.61 | -40.08 | -14.76 | -68.35 | -50.79 | -32.38 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 17.79 | 195.62 | 69.68 | 142.01 | 13.79 | 17.05 | 32.32 | 69.75 | ||
| ar | -40.49 | 92.62 | -91.01 | -47.79 | -85.73 | -73.79 | -33.37 | -39.94 | |
| de | -48.82 | 60.18 | -84.11 | -27.92 | -79.20 | -58.86 | -55.94 | -42.10 | |
| en | -31.83 | 51.71 | -80.42 | -25.18 | -55.99 | -55.87 | -50.75 | -35.47 | |
| es | -51.49 | 63.91 | -79.88 | 19.69 | -46.90 | -72.75 | -25.39 | -27.54 | |
| hi | -73.29 | 30.61 | -64.19 | 17.28 | -105.15 | -30.51 | -53.38 | -39.80 | |
| ru | -85.55 | 66.24 | -97.59 | 32.90 | -40.96 | -28.45 | -26.81 | -25.75 | |
| zh | -67.22 | 69.90 | -28.78 | 21.68 | -38.72 | -29.45 | -40.33 | -16.13 | |
| - () | AVG | -56.96 | 62.17 | -75.14 | -1.33 | -64.66 | -49.95 | -40.85 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 06.69 | 2259.35 | 04.01 | 16.44 | 12.06 | 564.88 | 04.82 | 409.75 | ||
| ar | -04.23 | 2250.85 | -08.09 | 08.15 | -01.69 | 552.11 | -03.24 | 399.12 | |
| de | -83.18 | 2116.93 | -76.53 | -88.91 | -382.00 | 479.28 | -67.29 | 271.18 | |
| en | 01.22 | 2255.43 | -05.20 | 10.95 | 04.74 | 557.96 | -01.99 | 403.30 | |
| es | -56.11 | 2220.91 | -08.01 | 11.14 | -11.81 | 526.42 | -33.18 | 378.48 | |
| hi | -04.21 | 2249.98 | -03.23 | 13.54 | 07.91 | 559.85 | -04.38 | 402.78 | |
| ru | -391.78 | 2187.84 | -260.01 | -34.54 | -445.47 | 358.98 | -107.10 | 186.85 | |
| zh | 02.82 | 2255.48 | 00.28 | 12.68 | 07.65 | 560.67 | 00.90 | 405.78 | |
| - () | AVG | -76.50 | 2219.63 | -51.54 | -9.57 | -117.24 | 513.61 | -30.90 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 23.57 | 114.45 | 44.15 | 145.82 | 159.41 | 314.27 | 187.03 | 141.24 | ||
| ar | 13.30 | 104.03 | 34.39 | 135.05 | 152.16 | 302.90 | 177.97 | 131.40 | |
| de | 14.90 | 84.26 | 06.48 | 133.17 | 75.15 | 218.62 | 149.80 | 97.48 | |
| en | 17.24 | 108.49 | 40.33 | 137.65 | 153.79 | 307.12 | 179.98 | 134.94 | |
| es | -06.20 | 108.73 | -143.29 | 142.12 | 151.53 | 29.67 | 184.90 | 66.78 | |
| hi | 13.55 | 103.51 | 35.75 | 136.49 | 156.85 | 310.41 | 177.81 | 133.48 | |
| ru | -77.36 | -199.05 | -10.32 | 128.23 | 28.62 | -388.32 | 172.26 | -49.42 | |
| zh | 19.55 | 111.16 | 40.13 | 142.39 | 152.46 | 310.17 | 183.87 | 137.10 | |
| - () | AVG | -0.72 | 60.16 | 0.50 | 136.44 | 124.37 | 155.80 | 175.23 |
| am | ar | de | en | es | hi | ru | AVG | ||
| 25.56 | 49.35 | 37.93 | 126.36 | 136.29 | 24.47 | 145.75 | 77.96 | ||
| ar | 15.81 | 35.69 | 28.49 | 116.88 | 128.85 | 13.48 | 137.82 | 68.15 | |
| de | 17.69 | -23.17 | -07.59 | 120.28 | 128.74 | -15.38 | 129.21 | 49.97 | |
| en | 20.08 | 44.56 | 33.89 | 118.54 | 130.26 | 17.29 | 138.87 | 71.93 | |
| es | 21.65 | 40.64 | 07.44 | 115.52 | 131.65 | -01.93 | 125.84 | 62.97 | |
| hi | 17.12 | 37.76 | 25.26 | 118.70 | 133.48 | 20.43 | 134.71 | 69.64 | |
| ru | -98.82 | -43.55 | -123.96 | 116.03 | 71.87 | -319.17 | 134.09 | -37.64 | |
| zh | 22.05 | 45.86 | 34.25 | 122.96 | 132.43 | 20.53 | 142.10 | 74.31 | |
| - () | AVG | 2.22 | 19.68 | -0.32 | 118.42 | 122.47 | -37.82 | 134.66 |